What is the equivalent resistance of these resistors? (c) 2.0 Ω 1.012 2.0 Ω Select one: A. 1.5Ω B. 0.2 Ω C.0.5 Ω D. 0.8 Ω E.5Ω
Answer #1

Let 'R' be the equivalent resistance

(1/R) = (1/R1) + (1/R2) + (1/R3)

=> (1/R) = (1/2) + (1/1) + (1/2)

=> (1/R) = 2

=> R = 0.5 Ω

option (c) is the correct answer